# thanks: These authors contribute equallythanks: These authors contribute equally

### v1

@@ Acknowledgements.
The results incorporating both $T_{1}$ and $T_{2}$ relaxation (combined amplitude and phase damping) indicate that even in the presence of strong decoherence, the quantum reservoir remains capable of nontrivial information processing. In the case of only amplitude damping, the effective dephasing rate is reduced to $(2T_{1})^{-1}$. Consequently, the optimal total STM capacity is reached at larger $\tau$ compared to the combined $T_{1}$ and $T_{2}$ case, as $T_{1}$ is roughly one order of magnitude larger than $T_{2}$ (see TABLE 3). Although one might expect that omitting $T_{2}$ dephasing would better preserve information encoded in coherence, simulations show no significant improvement in STM capacity. This is likely because a slow relaxation ($\sim 10^{-1}~{\rm Hz}$) tends to take effect over a long evolution ($\sim 10^{0}~{\rm s}$), during which the high-frequency Hamiltonian ($\sim{10^{2}}~{\rm Hz}$) causes the inputs to be largely mixed in a nonlinear manner. In this context, the additional dephasing appears beneficial for managing memory. Overall, improved performance is anticipated by jointly tuning both the relaxation rate and the input interval to suit specific task requirements Sannia et al. (2024).

### v2

@@ Appendix I Effects of $T_{1}$ and $T_{2}$ relaxation on the fading memory in quantum reservoir computing
The results incorporating both $T_{1}$ and $T_{2}$ relaxation (combined amplitude and phase damping) indicate that even in the presence of strong decoherence, the quantum reservoir remains capable of nontrivial information processing. In the case of only amplitude damping, the effective dephasing rate is reduced to $(2T_{1})^{-1}$. Consequently, the optimal total STM capacity is reached at a larger $\tau$ compared to the combined $T_{1}$ and $T_{2}$ case, as $T_{1}$ is roughly one order of magnitude larger than $T_{2}$ (see TABLE 3). Although one might expect that omitting $T_{2}$ dephasing would better preserve information encoded in coherence, simulations show no significant improvement in STM capacity. This is likely because a slow relaxation ($\sim 10^{-1}~{\rm Hz}$) tends to take effect over a long evolution ($\sim 10^{0}~{\rm s}$), during which the high-frequency Hamiltonian ($\sim{10^{2}}~{\rm Hz}$) causes the inputs to be largely mixed in a nonlinear manner. In this context, the additional dephasing appears beneficial for managing memory. Overall, improved performance is anticipated by jointly tuning both the relaxation rate and the input interval to suit specific task requirements [73].